12. Show that T:R → R given by T(2) = + 1 is an isomorphism.
Answer #1

We define Tirth by T(X) = x+1 - We take X, Y ER such that T(X) = T(X2) Əx,+1 = x3 + 1 it is one-to-one Assume any YER *4-1ER